# CoherenDream: Boosting Holistic Text Coherence in 3D Generation via Multimodal Large Language Models Feedback

### v1

@@ 4.3 Framework of CoherenDream
LLM-layout Initialization. As discussed in Sec. 4.1, the rendered images x of 3D representations significantly influence the sampled image distribution from the diffusion model. The usage of a unit sphere for initialization provides less information condition, leading to the traption of local minimum during the score distillation optimization. It also leads to textual inconsistency. To address this issue, we introduce an LLM-generated layout to warm up CoherenDream generation, which we refer to as LLM-Layout. As illustrated in the “Layout Condition Generation” part of Fig. 3, LLM-Layout consists of a set of cubes in normalized space. The collection of LLM-Layouts mirrors the layout condition generation process detailed in Sec. 4.2.2. Specifically, we fit the density network of NeRF to align with the occupancy defined by the LLM-Layout. Since the LLM-Layout only provides fundamental placement information, we do not wish to constrain the 3D representation’s flexibility regarding shape. To balance this, we implement a decay in importance near the surface, allowing for more freedom in shape representation. The surface weight decay binary cross entropy loss function can be expressed as:
@@
[…]
MVDream [34] 39.0 28.5 57min
ProlificDreamer [41] 45.7 25.8 240min
RichDreamer [31] 34.8 22.0 70min
CoherenDream 39.4 29.1 60min
Table 1: Comparison on T3Bench multiple objects set.
Method CLIP Score $\uparrow$ TIFA Score $\uparrow$ VQAScore $\uparrow$
MVDream [34] 30.5 77.4 0.73
DreamView [44] 31.1 77.9 0.71
JointDreamer [15] 30.8 73.7 0.71
CoherenDream 31.4 81.4 0.79
@@
Figure 5: Comparing TCSD with Original SDS and GraphDreamer. Original SDS exhibits bias accumulation issues, resulting in object omissions. GraphDreamer, on the other hand, may encounter unnatural decomposition, leading to objects being fully obscured by others. In contrast, TCSD leverages a dynamic MLLM assessor to produce coherent results in a holistic 3D space.
5 Experiments
In this part, we present the text-to-3D generation results of CoherenDream with qualitative and quantitative evaluations, illustrating state-of-the-art performance. We also make an ablation analysis of the proposed TCSD. More details can be found in the Appendix.
5.1 Textual Consistent 3D Generation
Qualitative Comparison.
Figure 4 shows the comparison with several representative baselines, results produced by official codes. (i) MVDream [34]: While MVDream finetunes multi-view diffusion model on 3D dataset demonstrating heavy textual inconsistency with multi-object prompts, our CoherenDream triggers generalization of multi-view diffusion model guiding by MLLMs feedback. (ii) DreamView [44]: DreamView finetunes on 3D dataset leading to unrealistic texture and implausible interaction (e.g., the mixture of “bear” and “toy car”). Additionally, it need view-specific prompts, which cost human labor. (iii) JointDreamer [15]: JointDreamer maintains the generalization of original 2D diffusion, but it cannot break through the inherent limitations of diffusion model. It demonstrates object omissions (“child” in the third line) and insufficient semantic interaction understanding (targets for “beside” but shows “on” in the second row). In comparison, our Conherent can produce more text-faithful results. (iv) GALA3D [46]: GALA3D utilizes a compositional generative pipeline, optimizing each object separately. However, the non-overlapping layout leads to unnatural object placement, like a “party hat” above the “koala”. (v) GraphDreamer [9]: GraphDreamer advances compositional optimization through scene graph decomposition. However, it tends to have insufficient optimization and unnatural object reorganization. For example, faceless cat and incorrect understanding to “leans” in the last row. Note that the same layout and scene graph are used for GALA3D, GraphDreamer, and CoherenDream to ensure a fair comparison.
Figure 6: Incremental ablations on techniques in CoherenDream framework, which contributes to enhancing text alignment.
Figure 7: Ablation Study on Guidance Tasks with TCSD. Scene Graph Generation task provides valuable global semantic guidance, resulting in coherent geometric structures, although it may overlook some text-specific details. The Multi-label Classification task ensures correspondence between objects and text; however, it lacks the incorporation of attribution and relational information, leading to unrealistic structures. CoherenDream integrates these three guidance tasks to enhance both text alignment and geometric fidelity in 3D results.
Quantitative Comparison.
@@ 5.1 Textual Consistent 3D Generation
We perform quantitative evaluations on two benchmarks: the multi-object set from T3Bench [10] and a curated 45-prompt subset of TIFA v1.0 [13]. For T3Bench, we employ two complementary metrics: ImageReward [43] for assessing 3D asset quality and GPT-4 [42] for text-to-3D alignment verification. As shown in Table 2, our CoherenDream achieves comparable mesh and texture quality to the baseline MVDream [34], while demonstrating superior performance in alignment accuracy. For the TIFA subset, we adopt a multi-metric evaluation protocol to evaluate textual consistency: CLIP score [11], TIFA score [13] and VQAScore [23].To ensure comprehensive viewpoint coverage, we uniformly sample 10 azimuth angles around each 3D asset. For VQAScore, we report the maximum score across all viewpoints; For TIFA Score, we compute the intersection of correct answers across all viewpoints. For Clip Score, we adopt the CLIP ViT-B/32 as the feature extractor and calculate the average score across views. Quantitative results in Table 2 demonstrate the consistent superiority of CoherenDream in text-aligned 3D generation across all metrics. Specifically, CoherenDream achieves an improvement of the TIFA score by 4.0% over MVDream, demonstrating its superior corresponding to textual description benefiting from effective feedback from MLLM. Detailed examples of TIFA’s question-answering pairs can be found in the Appendix.
5.2 Ablation Study
More Comparison with GraphDreamer. As show in Fig. 5, we conduct further comparison with GraphDreamer [9]. For a fair comparison, our TCSD is only equiped with scene graph generation task and no LLM-layout initialization. The results demonstrate that GraphDreamer encounters unnatural reorganization, leading to objects being fully obscured by others. In contrast, our TCSD optimizes a holistic 3D representation and receives informative feedback from MLLMs.
@@ 5.2 Ablation Study
Ablation on Textual Coherent Score Distillation. To verify the effectiveness of text-to-3D alignment achieved by SCSD, we distill from the multi-view diffusion model used in MVDream [34], which is known to overfit on 3D datasets. And we only utilize the original LLaVA-ov-0.5b with scene graph generation task as guidance. The primary distinction between the original SDS and SCSD, as illustrated in Fig. 1, is the introduction of an optimization direction $\delta$, which aims for ideal text alignment distribution defined in Eq. 6. As discussed in Sec. 4.1, the original SDS lacks the understanding and reasoning capabilities necessary for proper alignment, leading to significant misalignment issues. The incorporation of $\delta$ serves to mitigate these issues. The results presented in Fig. 5 validate our claims regarding the effectiveness of this approach.
@@
Ablation on LLM-Layout Initialization and 3DLLaVA-CRITIC. We conduct incremental ablations on techniques in CoherenDream, focusing on LLM-Layout Initialization and the fine-tuned 3DLLaVA-CRITIC. As shown in Fig. 6, LLM-Layout Initialization enhances the warm-up process for the 3D representation, effectively activating semantic information within the diffusion model and preventing the optimization process from collapsing into local minima. Compared to the scenario without LLM-Layout, this technique helps the model focus on the primary objects and establishes reasonable relative scale relationships. Moreover, the proposed LLM-Layout Initialization strategy does not constrain the creativity of 3D presentations. However, due to the domain gap between natural images and the sampled image distribution from the diffusion model, the semantic feedback provided by the original MLLM does not accurately correspond to object attributes and secondary objects, such as ”a parrot is two” and ”tuxedo” in Fig. 6 are overlooked. In contrast, our fine-tuned 3DLLaVA-CRITIC offers improved text-to-3D understanding feedback, resulting in more text-faithful 3D content.
Ablation on Guidance Tasks. To discuss the effectiveness of the decomposed guidance tasks, we conduct studies based on the overall framework of CoherenDream, replacing the combined guidance tasks with individual ones. As shown in Fig. 7, the scene graph generation task provides global semantic feedback and ensures geometric fidelity. In contrast, the multi-label classification task focuses on the objects present in the image distribution but ignores their attributes and relationships, which can lead to unrealistic placements. The view classification task enhances 3D consistency but it is similar with multi-label classification, ignoring object’s attribute. Additionally, view classification requires encoding the camera pose as $\textbf{T}_{vc}$, resulting in slower training speeds. Our proposed CoherenDream integrates these three guidance tasks to enhance both text alignment and geometric fidelity in 3D contents.
6 Conclusion

### v3

@@ Introduction
LLM-layout Initialization.
As discussed in Textual Coherent Score Distillation, the rendered images x of 3D representations significantly influence the sampled image distribution from the diffusion model. The usage of a unit sphere for initialization provides less information condition, leading to the traption of local minimum during the score distillation optimization. It also leads to textual inconsistency. To address this issue, we introduce an LLM-generated layout to warm up CoherenDream generation, which we refer to as LLM-Layout. As illustrated in the “Layout Condition Generation” part of Fig. 3, LLM-Layout consists of a set of cubes in normalized space. The collection of LLM-Layouts mirrors the layout condition generation process detailed in 3DLLaVA-CRITIC. Specifically, we fit the density network of NeRF to align with the occupancy defined by the LLM-Layout. Since the LLM-Layout only provides fundamental placement information, we do not wish to constrain the 3D representation’s flexibility regarding shape. To balance this, we implement a decay in importance near the surface, allowing for more freedom in shape representation. The surface weight decay binary cross-entropy loss function is:
@@
Experiments
We present the text-to-3D generation results of CoherenDream with qualitative and quantitative evaluations, illustrating state-of-the-art performance. We also make an ablation analysis of the proposed TCSD and 3DLLaVA-CRITIC. More details and experiments can be found in the Appendix.
Textual Consistent 3D Generation
Qualitative Comparison.
Figure 4 shows the comparison with several representative baselines, results produced by official codes. (i) MVDream (Shi et al. 2024): While MVDream finetunes multi-view diffusion model on 3D dataset demonstrating heavy textual inconsistency with multi-object prompts, our CoherenDream triggers generalization of multi-view diffusion model guiding by MLLMs feedback. (ii) DreamView (Yan et al. 2025): DreamView finetunes on 3D dataset leading to unrealistic texture and implausible interaction (e.g., the mixture of “bear” and “toy car”). Additionally, it need view-specific prompts, which cost human labor. (iii) JointDreamer (Jiang et al. 2025): JointDreamer maintains the generalization of original 2D diffusion, but it cannot break through the inherent limitations of diffusion model. It demonstrates object omissions (“child” in the third line) and insufficient semantic interaction understanding (targets for “beside” but shows “on” in the second row). In comparison, our Conherent can produce more text-faithful results.
Method TIFA Score $\uparrow$ VQAScore $\uparrow$ CLIP Score $\uparrow$
MVDream (Shi et al. 2024) 77.4 0.73 30.5
DreamView (Yan et al. 2025) 77.9 0.71 31.1
JointDreamer (Jiang et al. 2025) 73.7 0.71 30.8
CoherenDream 81.4 0.79 31.4
@@
Method $Acc_{\textbf{T}_{sg}}\uparrow$ $\textbf{T}_{ob}\uparrow$ $\textbf{T}_{view}\uparrow$
LlaVA-OV-0.5B (Li et al. 2024a) 0.21 0.34 0.72
3DLlava-CRITIC 0.76 0.89 0.87
Table 2: Quantitative evaluation for 3DLlava-CRITIC.
@@
Quantitative Comparison. We perform quantitative evaluations on a curated 45-prompt subset of TIFA v1.0 (Hu et al. 2023). We adopt a multi-metric evaluation protocol to evaluate textual consistency: TIFA score (Hu et al. 2023), VQAScore (Lin et al. 2024) and CLIP score (Hessel et al. 2021).To ensure comprehensive viewpoint coverage, we uniformly sample 10 azimuth angles around each 3D asset. For VQAScore, we report the maximum score across all viewpoints; For TIFA Score, we compute the intersection of correct answers across all viewpoints. For Clip Score, we adopt the CLIP ViT-B/32 as the feature extractor and calculate the average score across views. Quantitative results in Table 1 demonstrate the consistent superiority of CoherenDream in text-aligned 3D generation across all metrics. Specifically, CoherenDream achieves an improvement of the TIFA score by 4.0% over MVDream, demonstrating its superior corresponding to textual description benefiting from effective feedback from MLLM.
Ablation Study
Figure 5: Comparing TCSD with Original SDS. Original SDS exhibits bias accumulation issues resulting in object omissions, while TCSD leverages a dynamic MLLM assessor to produce coherent results in a holistic 3D space.
@@
Ablation on TCSD. To verify the effectiveness of text-to-3D alignment achieved by TCSD, we distill from the multi-view diffusion model used in MVDream (Shi et al. 2024), which is known to overfit on 3D datasets. And we only utilize the original LLaVA-ov-0.5b with scene graph generation task as guidance. The primary distinction between the original SDS and TCSD, is the introduction of an optimization direction $\delta$, which aims for ideal text alignment distribution defined in Eq. 6. Original SDS lacks the understanding and reasoning capabilities necessary for proper alignment, leading to significant misalignment issues. The incorporation of $\delta$ serves to mitigate these issues. The results presented in Fig. 5 validate our claims regarding the effectiveness of this approach.
Figure 6: Incremental ablations on techniques in CoherenDream framework, which enhances text alignment.
Figure 7: Compare with native 3D model in text-to-image-to-3D pipeline.
Ablation on LLM-Layout Initialization and 3DLLaVA-CRITIC.
@@
We conduct incremental ablations on techniques in CoherenDream, focusing on LLM-Layout Initialization and the fine-tuned 3DLLaVA-CRITIC. As shown in Fig. 6, LLM-Layout Initialization enhances the warm-up process for the 3D representation, effectively activating semantic information within the diffusion model and preventing the optimization process from collapsing into local minima. Compared to the scenario without LLM-Layout, this technique helps the model focus on the primary objects and establishes reasonable relative scale relationships. Moreover, the proposed LLM-Layout Initialization strategy does not constrain the creativity of 3D presentations.
However, the domain gap between real images and diffusion‐sampled views causes off‐the‐shelf MLLMs to miss fine attributes and objects (e.g., ignore “a parrot is two” or “tuxedo” as shown in Fig. 6. To address this, we fine‐tune 3DLLaVA-CRITIC, which delivers more accurate text‐to‐3D feedback and yields significantly more faithful 3D outputs. We also quantitatively compare 3DLLaVA-CRITIC against LLaVA-OV-0.5B in Table 5. In the absence of a ground-truth dataset, we construct 30 prompts for $\textbf{T}_{sg}$ and $\textbf{T}_{ob}$, and randomly choose 10 objects across 10 camera poses for $\textbf{T}_{view}$. For each validation sample, GPT-4o is given both models’ responses alongside the reference answer and tasked with scoring each as correct (1) or incorrect (0). The resulting average accuracies confirm the consistent gains achieved by our fine-tuning.
Compare with native 3D models.
Recently, native 3D generation models have attracted attention due to their rapid inference. However, constrained by the limited size and diversity of existing 3D datasets, these native models underperform optimization‐based approaches on complex or lengthy textual prompts. Fig. 7 presents a qualitative comparison between our method and several native pipelines, including Craftsman3D (Li et al. 2024c), Direct3D (Wu et al. 2024a), and Hunyuan3D-2 (Zhao et al. 2025). Most native models do not generate 3D geometry directly from text; instead, they employ a multistage text-image-textured-mesh pipeline. As a result, the final mesh quality depends heavily on the intermediate image and often omits fine details. Furthermore, errors introduced during text‐to‐image synthesis propagate through each stage, leading to significant misalignment between the user’s prompt and the native 3D output. These pipelines also tend to produce less realistic textures than SDS-based methods. Finally, although native models offer fast image‐to‐3D conversion, applying textures to an initially untextured mesh remains a computationally expensive step. In contrast, our CoherenDream maintains higher fidelity to textual prompts and produces more detailed, semantically consistent 3D content.
Conclusion
@@
Limitation. CoherenDream excels at holistic 3D generation and capturing intricate object interactions, but may underperform compositional methods in precise object arrangement and partial attribute control. But our methods can be integrated with compositional pipelines to harness complementary strengths.